# for GitHub Issue ReSolution

### v1

@@ Abstract
In software evolution, resolving the emergent issues within GitHub repositories is a complex challenge that involves not only the incorporation of new code but also the maintenance of existing functionalities. Large Language Models (LLMs) have shown promise in code generation and understanding but face difficulties in code change, particularly at the repository level. To overcome these challenges, we empirically study the reason why LLMs mostly fail to resolve GitHub issues and analyze some impact factors. Motivated by the empirical findings, we propose a novel LLM-based Multi-Agent framework for GitHub Issue reSolution, MAGIS, consisting of four kinds of agents customized for the software evolution: Manager, Repository Custodian, Developer, and Quality Assurance Engineer agents. This framework leverages the collaboration of various agents in the planning and coding process to unlock the potential of LLMs to resolve GitHub issues. In experiments, we employ the SWE-bench benchmark to compare MAGIS with popular LLMs, including GPT-3.5, GPT-4, and Claude-2. MAGIS can resolve 13.94% GitHub issues, which significantly outperforms the baselines. Specifically, MAGIS achieves an eight-fold increase in resolved ratio over the direct application of GPT-4, the based LLM of our method. We also analyze the factors for improving GitHub issue resolution rates, such as line location, task allocation, etc.
@@ 1 Introduction
In real-world software development, the software application is rarely set in stone. High-quality and popular software continually evolves to address emergent bugs or adapt to new requirements arising from shifts in user needs, software environments, and physical hardware. On platforms such as GitHub  [^1] https://github.com, issues typically signify the requirement for software evolution. However, addressing these issues poses significant challenges, as it requires implementing the code change across the entire repository and maintaining the existing functionality while integrating new capabilities. Consequently, resolving GitHub issues remains a significant challenge across academia and industry [15, 4].
@@
Large Language Models (LLMs) have demonstrated remarkable capabilities across a variety of tasks [6], including code generation and code understanding [40, 30]. Specifically, LLMs excel in generating function-level code, as evidenced by their performance on numerous benchmark datasets such as MBPP [2] and HumanEval [9]. Despite their success, LLMs remain challenged in tasks that require advanced code generation capabilities, such as the ClassEval benchmark [10]. Moreover, LLMs exhibit limitations in processing excessively long context inputs and are subject to constraints regarding their input context length [20]. This limitation is particularly evident in repository-level tasks, such as solving GitHub issues, where the context comprises the entire repository, thus imposing constraints on directly using the full repository as input to LLMs.
@@
To harness the full potential of LLMs, researchers [12, 26, 36] have designed LLM-based multi-agent systems. These approaches have significantly improved LLMs’ efficacy in code generation, enabling these systems to construct code repositories based on LLM. While these methods address the process of transitioning code repositories from inception to establishment, they rarely consider the handling of software evolution, e.g., resolving Github issues. For Github repositories, especially popular ones, a large number of commits are pushed every day. These commits derive from a spectrum of evolutionary requirements that span bug fixes, feature additions, performance enhancements, etc [32]. For open-source software, new requirements frequently emerge as issues in the project’s repository.
@@
To investigate the capability of LLMs in addressing GitHub issues, Jimenez et al. [15] developed a benchmark, namely SWE-bench, to conduct a comprehensive analysis of popular LLMs. The study reveals that LLMs fail to resolve over $5\%$ of instances, even when provided with file paths that require modifications. This significantly low rate of applying solutions and resolving issues underscores the limitations inherent in LLMs. The reasons behind the suboptimal performance of LLMs and strategies to harness their potential for GitHub issue resolution remain under-explored.
@@
In this study, we analyze the factors impacting the effectiveness of LLMs in resolving issues under two different settings (i.e., w/ Oracle and w/o Oracle). Furthermore, our empirical analysis has concluded a correlation between file location and line location within files, and performance in resolving GitHub issues. Based on these insights, we propose a novel LLM-based Multi-agent Framework, termed MAGIS, comprising four types of agents: Manager, Repository Custodian, Developer, and Quality Assurance (QA) Engineer. Our approach facilitates the resolution of GitHub issues through collaboration among agents, each fulfilling a unique role.
@@
In our experiment, we evaluate our framework on the SWE-bench, comparing its performance against existing popular LLMs, such as ChatGPT-3.5 [24], GPT-4 [25], and Claude-2 [1]. The results demonstrate that our framework, utilizing GPT-4 as its base model, significantly outperforms the baseline, and achieves an eight-fold performance gain compared to the direct application of GPT-4. Further analysis revealed additional factors, i.e., the planning of code change, line location within the code file, and code review process, that significantly influence the resolution rate.
@@
  • We conduct an empirical analysis of LLMs in resolving GitHub issues in two settings (i.e., w/ Oracle setting and w/o Oracle setting), and explore the correlation between line location, complexity of the code change, file location, and the success rate in resolving GitHub issues.
  • We propose a novel LLM-based multi-agent framework, MAGIS, to alleviate the limitations of existing LLMs on GitHub issue resolution. Both our designed four-type agents and their collaboration for planning and coding unlock LLMs’ potential on the repository-level coding task.
  • We compare our framework and other strong LLM competitors (i.e., GPT-3.5, GPT-4, and Claude-2) on the SWE-bench dataset. The results show MAGIS significantly outperforms these competitors. Further analysis is conducted and verifies the effectiveness and necessity of our framework design.
@@ 2 Empirical Study
2.1 RQ 1: What Impacts the Performance Under With-Oracle Setting?
The with-Oracle setting is a more straightforward setup for the GitHub issue resolution by supplying the specific files requiring modifications. Despite the observation that LLMs exhibit better performance in the with-Oracle setting compared to their performance without Oracle, the proportion of issues successfully resolved under the with-Oracle setting remains modest. Specifically, in the with-Oracle setting, GPT-4 achieved a resolved rate of only $1.74\%$ on the SWE-bench.
To delve into the factors impacting the efficacy within the with-Oracle setting, we examined failed instances’ generation. This scrutiny led to the identification of two factors that could influence the resolved rate: (1) Line location (location of the modified line in the code file), and (2) The complexity of the code change. This information is important and provided in the code change of each commit in the code repository. A typical code change includes multiple changed hunks and each hunk contains the line numbers targeted for modification and the specifics of the changed code at these locations. Both the line locations and the changed content make a repository evolve.
To validate this identification, we analyze the relation between the resolved rate and the two factors mentioned above.
2.1.1 Line Location
@@
To quantitatively analyze the accuracy of line localization, we use the line numbers’ range of the modified content in the reference code change as the basis assuming that the correct modification location of the code change is uniquely determined in most cases. By calculating the overlap ratio of the line number ranges of the generated and reference, we can estimate the accuracy of line localization in the generation process. Specifically, for each instance, the line number ranges of the code change in the reference $r$ is represented as a set of intervals ${\bm{L}}_{r}=\{[s_{0},e_{0}],...,[s_{n},e_{n}]\}$, while the line number range set of the generated code change $g$ is ${\bm{L}}_{g}=\{[s_{0}^{\prime},e_{0}^{\prime}],...,[s_{m}^{\prime},e_{m}^{\prime}]\}$, where $s$ and $e$ respectively represent the starting and ending line number of each modification hunk in the file, with $n$ hunks in the reference code change and $m$ in the generated one. The formula for calculating the overlap ratio is as follows:
Overlap $\text{Overlap Ratio}=\frac{\sum_{i=0}^{n}\sum_{j=0}^{m}\left|[s_{i},e_{i}]\cap[s_{j}^{\prime},e_{j}^{\prime}]\right|}{\sum_{i=0}^{n}(e_{i}-s_{i}+1)},~{}$ (1)
where $\left|[s_{i},e_{i}]\cap[s_{j}^{\prime},e_{j}^{\prime}]\right|$ denotes the size of the intersection between the intervals $[s_{i},e_{i}]$ and $[s_{j}^{\prime},e_{j}^{\prime}]$.
For $574$ instances in the SWE-bench test set that experiments GPT-4, the distribution of the overlap ratio between the results generated by three LLMs and the reference code change is shown in Figure 1 with the vertical axis representing the frequency of the range of line location overlap ratio for each group, and the horizontal axis representing the overlap ratio. From this, we observe that: (1) The distribution near the overlap ratio 00 (left side of the figure) is the highest for all three LLMs, indicating that in most cases, the content generated by these models has a very low overlap ratio with the reference in terms of line location. This means that these LLMs are most likely not able to accurately locate the lines of code that need to be modified in the process of generating the code change. (2) In the distribution near the line location overlap of $1$ (right side of the figure), the three models show a consistent ranking (i.e., Claude-2 $>$ GPT-4 $>$ GPT-3.5) and this ranking is also consistent with the proportion of instances solved by the three models. This phenomenon suggests that the performance of LLMs in generating the code change is probably related to their ability to locate code lines accurately.
Furthermore, we assess the relationship between the overlap ratio and the resolution of GitHub issues by calculating their correlation coefficient. Given that the distribution of these variables exhibits
Figure 1: The Comparison of Line Location Overlap Ratio between Three LLMs On SWE-bench.
@@
skewness, and the resolution result is binary (resolved or not), logistic regression is employed for the analysis across three LLMs. However, due to the limited number of successfully generated instances on GPT-4 and GPT-3.5, a statistically significant relationship (P-value < $0.05$) is detected solely in the generated result from Claude-2. This signifies that the correlation coefficient from Claude-2 is both statistically significant and meaningful. Specifically, with a positive coefficient, $0.5997$, on Claude-2, there is a substantial and positive relation between improvements in the overlap ratio and the probability of successfully resolving issues. Consequently, the line location emerges as an important factor for GitHub issue resolution.
2.1.2 Complexity of the Code Change
@@
The complexity of the code change is reflected in various indices, including the number of modified files, functions, hunks, and lines added or deleted. Firstly, we quantitatively assess the complexity by calculating the value of various indices corresponding to the reference code change. Secondly, the coefficient is calculated between the numbers in each index and the resolution of issues. Table 1 shows the correlation scores between six indices and the issue resolution, under the logistic regression.
@@
As shown in Table 1, all three LLMs, i.e., GPT-3.5, GPT-4, and Claude-2, demonstrate a statistically significant correlation with the issue resolution across several indices, i.e., p-value less than $0.05$ and marked by *. The correlation scores for the number of files and functions modified are notably negative for all models, indicating that an increase in these indices is associated with a decreasing likelihood of issue resolution. This suggests that the more complex the code change, as indicated by a higher number of files and functions modified, may hinder the issue resolution. Compared with GPT-3.5 and GPT-4, Claude-2 exhibits a different pattern, with much lower negative correlations for the number of files and functions, which indicates it is a more efficient approach to generate the code change for GitHub issue resolution. However, it also shows significant negative correlations across other indices such as the number of hunks, added lines of code (LoC), deleted LoC, and changed LoC. The analysis reveals a relationship between the complexity, as measured by several indices, and whether to successfully resolve the issues in software evolution. The negative correlations suggest that increased complexity, particularly in terms of the number of files and functions changed, tends to hinder issue resolution.
2.2 RQ 2: What Impacts the Performance Under Without-Oracle Setting?
The difference between the with-Oracle setting and the without-Oracle setting lies in the necessity for the latter to identify the modified files. Consequently, file locating emerges as a crucial factor influencing performance under the without-Oracle setting.
Jimenez et al. [15] employ the BM25 method [28] to retrieve relevant code files that are subsequently utilized as input to the LLM. After employing retrieval methods, it is necessary to select the top-$K$ files or truncate the content based on the maximum context length of the LLM. Incorporating more files can enhance recall scores. However, it also imposes significant demands on the capabilities of LLMs. As demonstrated by the study [15], Claude-2 exhibits a reduction at the resolved ratio as recall scores increase. This decline may be attributable to the inclusion of irrelevant files or the limited capacity of LLMs to process longer contexts effectively. Consequently, exploiting the performance of LLMs can be better achieved by striving for higher recall scores with a minimized set of files, thus suggesting a strategic balance between recall optimization and the number of chosen files.
@@ 3 Methodology
In this section, we elaborate on our LLM-based multi-agent framework, MAGIS, for GitHub issue resolution. The framework resolves each issue by applying the generated code change to the repository. As shown in Figure 2, we first provide the overview of our framework (§3.1), followed by the role design of each type agent and how these roles differ from human workflows (§3.2). Lastly, we detail the collaborative process among these roles within our framework (§3.3).
3.1 Overview
The architecture of our framework is illustrated in Figure 2, encompassing four key roles of agents working collaboratively in the workflow.
  • Manager. This role tasks with team assembly, meeting organization, and plan formulation.
  • Repository Custodian. The custodian is responsible for locating the relevant files in the repository according to the GitHub issue and recording the change of the repository.
  • Developer. This role participates in planning discussions and completes tasks from the manager.
  • Quality Assurance (QA) Engineer. The QA engineer reviews the code change from developers to ensure the quality of the whole repository.
Figure 2: Overview of Our Framework, MAGIS.
The collaborative process involves planning and coding. In the planning, a GitHub issue is assigned to the project manager and the repository custodian. The repository custodian identifies candidate files relevant to the issue for modification. With the issue description and a list of candidate files, the manager defines tasks and assembles a team, where each member is a developer specifically designed for the defined task. The manager holds a kick-off team meeting with developers and devises a plan. During the coding, developers undertake their assigned tasks from the manager, and the QA engineer reviews each code change. If a change fails to meet quality standards, the QA engineer provides feedback, prompting further revisions until the change satisfies quality criteria or a set iteration limit is reached.
@@ 3.2 Agent Role Design
Our workflow draws inspiration from the GitHub Flow  [^2] https://docs.github.com/en/get-started/using-github/github-flow, an effective human workflow paradigm, adopted by many software teams. Both the human workflow and our LLM-based agent framework prioritize collaboration among individuals with diverse skills. While the underlying principles are similar, there are notable differences. Accordingly, we have tailored the roles as follows:
@@
  • [Uncaptioned image] Repository Custodian. Considering extensive files in a repository, the custodian agent’s task is to locate files relevant to the issue. Different from humans, who can browse through the entire repository, the LLM-based agent faces challenges in browsing. Although LLMs, including GPT-4 [23], have extended context limits, their application is constrained in two aspects. First, it is a high computational cost to query the whole repository for each update while some repositories update frequently. Second, the performance of LLMs degrades when relevant information occurs in the middle of the long context input [19, 42].
  • [Uncaptioned image] Developer. In contrast to human developers, the developer agent can work continuously, and complete a task in minutes even seconds. Therefore, scheduling the agent to work in parallel is easier than scheduling humans as the latter requires considering other than work while the former only needs to ensure the tasks are independent. Additionally, although there are numerous developer agents capable of generating code [12, 26], their ability to modify existing code is not equally proficient. To address this issue, our framework decomposes the code modification process into atomic operations, which encompass the code generation. This approach enables developers to leverage the benefits of automatic code generation thereby generating applicable code changes.
  • [Uncaptioned image] QA Engineer. In software evolution, QA engineers, especially code reviewers, play a crucial role in maintaining software quality [21, 17]. Despite its critical role, code review practices are often undervalued or even overlooked [3]. Such neglect can hinder software development, illustrated by instances where developers may experience delays of up to 96 hours awaiting code review feedback [5]. To address this problem, our framework pairs each developer agent with a QA engineer agent, designed to offer task-specific, timely feedback. This personalized QA approach aims to boost the review process thereby better ensuring the software quality.

### v2

@@ Abstract
In software development, resolving the emergent issues within GitHub repositories is a complex challenge that involves not only the incorporation of new code but also the maintenance of existing code. Large Language Models (LLMs) have shown promise in code generation but face difficulties in resolving Github issues, particularly at the repository level. To overcome this challenge, we empirically study the reason why LLMs fail to resolve GitHub issues and analyze the major factors. Motivated by the empirical findings, we propose a novel LLM-based Multi-Agent framework for GitHub Issue reSolution, MAGIS, consisting of four agents customized for software evolution: Manager, Repository Custodian, Developer, and Quality Assurance Engineer agents. This framework leverages the collaboration of various agents in the planning and coding process to unlock the potential of LLMs to resolve GitHub issues. In experiments, we employ the SWE-bench benchmark to compare MAGIS with popular LLMs, including GPT-3.5, GPT-4, and Claude-2. MAGIS can resolve 13.94% GitHub issues, significantly outperforming the baselines. Specifically, MAGIS achieves an eight-fold increase in resolved ratio over the direct application of GPT-4, the advanced LLM.
@@ 1 Introduction
In real-world software development, the code repository for a project is rarely set in stone. High-quality and popular software always evolves to address emergent bugs or new requirements. On platforms such as GitHub [21], issues typically signify the requirement for software evolution. However, addressing these issues poses significant challenges, as it requires implementing the code change across the entire repository and maintaining the existing functionality while integrating new capabilities. For example, django, a framework for over $1.6$M projects has $34$K issues [19]. Consequently, resolving GitHub issues remains a significant challenge across academia and industry [27, 5].
@@
Large language models (LLMs) have demonstrated remarkable capabilities across a variety of tasks [8], including code generation and code understanding [65, 48]. Specifically, LLMs excel in generating function-level code, as evidenced by their performance on numerous benchmark datasets such as MBPP [2] and HumanEval [12]. Despite their success, LLMs remain challenged in tasks that require advanced code generation capabilities, such as class-level code generation [14]. Moreover, LLMs exhibit limitations in processing excessively long context inputs and are subject to constraints regarding their input context length [33]. This limitation is particularly evident in repository-level coding tasks, such as solving GitHub issues, where the context comprises the entire repository, thus imposing constraints on directly using the full repository as input to LLMs.
@@
To harness the full potential of LLMs, many LLM-based multi-agent systems are designed [23, 44, 53]. These methods have significantly improved LLMs’ efficacy in code generation, enabling these systems to construct code repositories based on LLM. While these methods address the process of transitioning code repositories from inception to establishment, they rarely consider the handling of software evolution, e.g., resolving GitHub issues. For GitHub repositories, especially the popular ones, a large number of commits are pushed every day. These commits derive from a spectrum of evolutionary requirements that span bug fixes, feature additions, performance enhancements, etc [50]. For open-source software, new requirements frequently emerge as issues in the project’s repository.
@@
Recently, Jimenez et al. [27] developed a benchmark, namely SWE-bench, to investigate the capability of popular LLMs in addressing GitHub issues. Their study reveals that LLMs fail to resolve over $95\%$ of instances, even when file paths that require modifications are provided. This significantly low rate underscores the importance of understanding the reasons behind their suboptimal performance.
@@
In this study, we analyze the factors impacting the effectiveness of LLMs in resolving GitHub issues. Furthermore, our empirical analysis has concluded a correlation between locating files/lines to be modified and the performance of resolving GitHub issues. Based on these insights, we propose a novel LLM-based multi-agent framework, termed MAGIS, comprising four types of agents: Manager, Repository Custodian, Developer, and Quality Assurance (QA) Engineer. Our approach facilitates the resolution of GitHub issues through collaboration among agents, each fulfilling a unique role: the Manager coordinates the entire process, the Repository Custodian enhances locating files, the Developer performs code changes after locating lines, and the QA Engineer reviews the code change.
@@
In our experiment, we evaluate our framework on SWE-bench and compare its performance against existing popular LLMs, such as ChatGPT-3.5 [37], GPT-4 [38], and Claude-2 [1]. The results demonstrate that our framework, utilizing GPT-4 as its base model, significantly outperforms baselines and achieves an eight-fold performance gain compared to the direct application of GPT-4. Further analysis reveals that additional factors, i.e., the planning of code change, locating lines within the code file, and code review process, can significantly influence the resolution rate.
@@
  • We conduct an empirical analysis of LLMs in resolving GitHub issues and explore the correlation between locating code file/line, complexity of the code change, and the success rate in resolution.
  • We propose a novel LLM-based multi-agennt framework, MAGIS, to alleviate the limitations of existing LLMs on GitHub issue resolution. Both our designed four-type agents and their collaboration for planning and coding unlock LLMs’ potential on the repository-level coding task.
  • We compare our framework and other strong LLM competitors (i.e., GPT-3.5, GPT-4, and Claude-2) on the SWE-bench dataset. The results show MAGIS significantly outperforms these competitors. Further analysis confirms the effectiveness and necessity of our framework design.
@@ 2 Empirical Study
SWE-bench [27] reveals the challenges LLMs face in addressing GitHub issue resolution. For example, in their evaluation, GPT-4 can only resolve less than 2% issues of the test set. Conversely, in tasks like function-level code generation, LLMs exhibit superior performance (e.g., GPT-4 gets the score of $67.0$ on HumanEval [36]). Given the complexity of GitHub issue resolution akin to repository-level coding, we aim to investigate Why the Performance of Directly Using LLMs to Resolve GitHub Issue is Limited? (RQ 1). We answer this RQ from the following three aspects:
Locating the Files to be Modified.
GitHub issue resolution is a repository-level coding task, distinguishing it from file-level coding tasks primarily in the challenge of locating the files requiring modification. Jimenez et al. [27] employ the BM25 method [46] to retrieve relevant code files that are subsequently utilized as input to the LLM. After employing retrieval methods, it is necessary to select the top-$K$ files or truncate the content based on the maximum context length of the LLM. Incorporating more files can enhance recall scores. However, it also imposes significant demands on the capabilities of LLMs. As demonstrated by the study [27], Claude-2 exhibits a decrease in the resolved ratio (from 1.96 to 1.22) as recall scores increase (from 29.58 to 51.06). This decline may be attributed to the inclusion of irrelevant files or the limited capacity of LLMs to process longer contexts effectively. Consequently, optimizing the performance of LLMs can be better achieved by striving for higher recall scores with a minimized set of files, thus suggesting a strategic balance between recall optimization and the number of chosen files.
Locating the Lines to be Modified.
@@
Beyond the impact of file locating, we delve into the generation of failed instances when the correct modified files were provided. A typical code change consists of multiple hunks, each specifying the line numbers targeted for modification and detailing the changes made at these locations. To quantitatively analyze the accuracy of line localization, we use the line numbers’ range of the modified content in the reference code change as the basis assuming that the correct modification location of the code change is uniquely determined in most cases. By calculating the coverage ratio of the line number ranges of the generated and reference, we can estimate the accuracy of line localization in the generation process, i.e.,
Coverage $\text{Coverage Ratio}=\frac{\sum_{i=0}^{n}\sum_{j=0}^{m}\left|[s_{i},e_{i}]\cap[s_{j}^{\prime},e_{j}^{\prime}]\right|}{\sum_{i=0}^{n}(e_{i}-s_{i}+1)},~{}$ (1)
where the numerator is the length of the intersection of modified lines between the reference divided into $n$ hunks and the generation divided into $m$ hunks, and the denominator is the number of modified lines in the reference. More details about Equation 1 can be found in Appendix A.1.
Figure 1: The comparison of line locating coverage ratio between three LLMs. The vertical axis representing the frequency of the range of line locating coverage ratio for each group, and the horizontal axis representing the coverage ratio.
For $574$ instances in the SWE-bench that experiments GPT-4 [27], the distribution of the coverage ratio between the results generated by three LLMs and the reference is shown in Fig. 1. From this, we observe that the performance of LLMs in generating the code change is probably related to their ability to locate code lines accurately (Detailed explanation can be found in Appendix A.2).
@@
Furthermore, we assess the relationship between the coverage ratio and the issue resolution by calculating their correlation coefficient. Given that the distribution of these variables exhibits skewness, and the resolution result is binary (resolved or not), logistic regression is employed for the analysis across three LLMs. However, due to the limited number of successfully generated instances on GPT-4 and GPT-3.5, a statistically significant relationship is only detected in the result generated by Claude-2. The result, i.e., P-value < $0.05$, shows statistical significance. Specifically, with a coefficient, $0.5997$, on Claude-2, there is a substantial and positive relation between improvements in the coverage ratio and the probability of successfully resolving issues, which demonstrates that locating lines is a key factor for GitHub issue resolution.
Complexity of the Code Changes.
@@
The complexity of the code change is reflected in various indices: the number of modified files, functions, hunks, and lines added or deleted. Firstly, we quantitatively assess the complexity by calculating the value of various indices corresponding to the reference code change. Secondly, the coefficient is calculated between the numbers in each index and the issue resolution. Tab. 1 shows the correlation scores under the logistic regression.
@@
As shown in Tab. 1, all three LLMs demonstrate a statistically significant correlation with the issue resolution across several indices. The correlation scores for the number of files and functions modified are notably negative for all models, indicating that an increase in these indices is associated with a decreasing likelihood of issue resolution. This suggests that the more complex the code change, as indicated by a higher number of files and functions modified, may hinder the issue resolution. More analysis can be found in Appendix A.3. The analysis reveals a relationship between the complexity, as measured by several indices, and whether to successfully resolve the issues in software evolution. The negative correlations suggest that increased complexity, particularly in terms of the number of files and functions changed, tends to hinder issue resolution.
@@ 3 Methodology
Based on the empirical study identifying key factors affecting LLMs’ issue resolution, we design the framework illustrated in Fig. 2. This framework aims to mitigate negative impacts by transforming the complex task of GitHub issue resolution into a collaborative effort. It incorporates four key roles for LLM-based agents working collaboratively in the workflow: ① Manager: this role tasks with team assembly, meeting organization, and plan formulation. ② Repository Custodian: it is responsible for locating the relevant files in the repository acording to the GitHub issue and recording the change of the repository. ③ Developer: this role participates in planning discussions and completes tasks from the Manager. ④ Quality Assurance (QA) Engineer: it reviews the code change from Developers to ensure the quality of the whole repository.
Figure 2: Overview of our framework, MAGIS. The detailed version can be found in Fig. 14.
The collaborative process involves planning and coding. In the planning, an issue is assigned to the Manager and the Repository Custodian. The custodian identifies candidate files relevant to the issue for modification. With the issue description and a list of candidate files, the Manager defines tasks and assembles a team, where each member is a Developer specifically designed for the defined task. The Manager holds a kick-off meeting with Developers and devises a plan. During coding, Developers undertake their assigned tasks from the Manager, and the QA Engineer reviews each code change. If a change fails to meet quality standards, the QA Engineer provides feedback, prompting further revisions until the QA Engineer approves or a set iteration limit is reached.
@@ 3.1 Agent Role Design
Our workflow draws inspiration from the GitHub Flow[22], an effective human workflow paradigm adopted by many software teams. Both the human workflow and our LLM-based agent framework prioritize collaboration among individuals with diverse skills. While the underlying principles are similar, there are notable differences. Accordingly, we have tailored the roles as follows:
@@
  • [Uncaptioned image] Repository Custodian. Considering extensive files in a repository, the custodian agent’s task is to locate files relevant to the issue. Unlike humans, who can browse through the entire repository, the LLM-based agent faces challenges in browsing. Although LLMs have extended context limits, their application is constrained in two aspects. First, it is a high computational cost to query each file in an entire repository for each update, particularly when some repositories update frequently. Second, the performance of LLMs degrades when the context input is long [31, 33, 68].
  • [Uncaptioned image] Developer. Compared to human developers, the Developer agent can work continuously and efficiently. Therefore, scheduling the agent to work in parallel is easier than scheduling humans who require considering factors beyond the task. Additionally, although numerous developer agents are capable of generating code [23, 44], their ability to modify existing code is not equally proficient. To address this issue, our framework decomposes the code modification process into sub-operations including code generation. This approach enables Developers to leverage the benefits of automatic code generation thereby producing applicable code changes.
  • [Uncaptioned image] QA Engineer. In software evolution, QA Engineers play a crucial role in maintaining software quality through code review [34, 30]. Despite their importance, code review practices are often undervalued or even overlooked [4]. Such neglect can hinder software development, illustrated by instances where developers may experience delays of up to 96 hours awaiting code review feedback [6]. To address this problem, our framework pairs each Developer agent with a QA Engineer agent, designed to offer task-specific, timely feedback. This personalized QA approach aims to boost the review process thereby better ensuring the software quality.